find the slope-intercept form of the line whos slope is 4 and passes through the point ( -9, 11)
Answer by mananth(16946)

use the formula
y-y1=m(x-x1)
4 = m (x1,y1) (-9,11)
y-11=4(x-(-9))
y-11=4x+36
y=4x+47